The trouble that I'm having is finding the exact answer this is how far I get then get stuck. 20 7/9 - 4 5/6 = 20 7/9 =21 7/9 and 4 5/6 =5 giving me 16 then to get the exact now i get stuck. I know how to get the 7/45 and 5/45 /
Answer by Alan3354(69443)

20 7/9 - 4 5/6
---------------------
= 187/9 - 29/6
Find a common denominator. The product always works, 6*9 = 54
= 6*187/54 - 9*29/54
= 1122/54 - 261/54
= 861/54
= 287/18